OpenStudy (anonymous):

Find the exact value of the trigonometric function at the given real number? - Sin (21pi/2) - Cos (21pi/2) - Cot (21pi/2)

OpenStudy (anonymous):

\[\sin(21\pi /2 ) = (21\pi/2 )(180/\pi) = \sin(1890) = 1\]\[\cos(21\pi / 2) = (21\pi/2)(180/\pi) = \cos(1890) = 0\]\[\cot(1890) = 1/\tan(1890) =\cos(1890)/\sin(1890) = 0\] All you have to do is convert the angle from radians to degree by multiplying the given angle with: \[180/\pi\] but if you want to convert from degree to radians then multiply by: \[\pi/180\] ^_^